OpenStudy (anonymous):

PLEASE help :) Let f(x)=4x+7. Find f(x+h)−f(x) all divided by (h)

OpenStudy (mertsj):

\[\frac{4(x+h)+7-(4x+7)}{h}=\frac{4x+4h+7-4x-7}{h}=\frac{4h}{h}=4\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (mertsj):

yw